OpenStudy (anonymous):

Please and Thank you! What is the simplified form of x-2/x^2+x-6 divided by x^2+5x+4/x+4 ?

OpenStudy (anonymous):

\[\frac{x-2}{x^2+x-6}\times \frac{x+4}{x^2+5x+4}\] is a start i.e. multiply by the reciprocal now factor and cancel as much as you can

OpenStudy (anonymous):

\[\frac{x-2}{x^2+x-6}\times \frac{x+4}{x^2+5x+4}\] \[=\frac{x-2}{(x+3)(x-2)}\times \frac{x+4}{(x+1)(x+4)}\] and then cancel away

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

don't forget to leave a \(1\) in the numerator yw

OpenStudy (anonymous):

1/(x+3)(x+1)

OpenStudy (anonymous):

yes, that works

OpenStudy (anonymous):

unfortunately problems like this are very misleading it says "multiply" or "divide" but it really means "factor and cancel" leading people to believe you can always cancel stuff, which is in fact rarely the case

OpenStudy (anonymous):

I see.. urgh Algebra-_-